Prev Next

SW&IT Meeting Minutes 2021-11-29

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20211115
LR20211115 - Bec, Matthieu 64.0
LR20211115 - Mechtley, Brandon 64.0
LR20211115 - Eychaner, Glenn 40.0
LR20211115 - Edmeades, David 64.0
LR20211115 - Grenz, Paul 12.0
LR20211115 - Hooper, Stephen 0.0
LR20211115 - Zhang, Yang 40.0
LR20211115 - Zhang, Xianyu 40.0

News from management

  • Reminder to file UAccess time when taking vacation

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • Performance Appraisals
    • shark-nir PAE
  • This week
    • Performance Appraisals
    • git/github "policy" (= philosophy slides) for new staff
    • shark-nir PAE ctd.
    • LBTplot conversion to DMS (ctd.)

YZ

  • Previous:
    • Vacation + Holidays
  • Next:
    • Modify and work on completing MODS WCS program (and probably schedule experiment with Olga).
    • Work on to finish the LBC bias level adjustment wiki page.
    • Resume TMS work and discussions.
    • New image notification tool according to archive meeting plan.
    • Argos swingarm/OSS test/debug/patch?
    • Other potential misc works on LBC/MODS if have time.

PSG

  • Last week
    • Proxy work, could not get testing time.
    • Looking at GCS code.
  • This week
    • Possibly have testing time on telescope
    • Make correction to GCS to fix error message.

DE

  • Last week
    • Set up new HA Synology system with 18TB drives, has been burning in. Looks great!
    • New Autodesk Vault 2020 running on Windows Server 2019, which is off-label but looks like it's working.
      • Plans to meet with Ketiv this week to refine install and plan transition
    • Worked on laptop spec for JW, he took the weekend to think about it
    • AC moving office, discovered inactive Ethernet ports. I opened a ticket with NetOps.
  • This week
    • Multiple drive issues on old Synology
    • Ketiv meeting
    • Test Vault backup script on new install; hopefully they didn't break anything fundamental
    • VM backup server
    • Still have Windows license key problems, but it's not affecting anything... yet. Will see if I can get answers.

GE

  • Last week
    • Digging through the 550,000 lines of code of the LUCI system.
    • Continued working on the RTD issues.
  • This week
    • Digging through the 550,000 lines of code of the LUCI system.
    • Create a pull request and merge the tested RTD into the LBTO GitHub.
    • Create a release version of the new LUCI software and place it in stow on the observer machines.
    • Continue working on the RTD issues.
    • Examine the other LUCI GUIs to check for GUI modifications from threads other than the Event Dispatch Thread.

SH

  • VMware mountain, PXE UEFI no go.
  • Saturday computer questions (Thanksgiving weekend!)
  • ROFIN battery.
  • Room 507 check
next:
  • broken monitors (talk to LeRoy Wednesday)
  • VMware (install, e-mail Arto, etc).
  • Quotes for switches, purchase second-hand switches
  • Wiki on NFS (Tucson Synology).
  • ROFIN battery?
Topic revision: r9 - 29 Nov 2021, StephenHooper
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback